OpenStudy (anonymous):

What is 2/3 x 5/6 x 14

OpenStudy (anonymous):

7.77

OpenStudy (anonymous):

2/3 *5/6 _______ 10/18 /2/2 = 5/9 5/9 * 14= 5/9 * 14/1 = 70/9 = 7-6/9 6/9/3 = 2/3 final answer 7-2/3 or in decimal form 7.77777777778